# The Dynamic Net Architecture: Learning Robust and Holistic Visual Representations Through Self-Organizing Networks

### v1

@@ 4 Computational Framework
Figure 2: The dynamics of a single area within the proposed DNA: The input image is fed into stage $S1$ to obtain feature activations. The feature activations are then processed by stage $S2$, together with activations from neurons within the same layer. In the upper part A, two distinct matrices are used to model the forward connections from $S1$ to $S2$ and the lateral connections within $S2$. In the lower part B, the matrices and input signals are concatenated, resulting in one bigger weight matrix.
We here restrict ourselves to modeling this functionality within a single cortical area, although the significance of coherent nets in primary sensory areas can be fully appreciated only in the context of different interacting areas.
A single area of a DNA model comprising net fragments is shown in Figure 2(A). Our system is based on an input layer representing observed images by simple feature detectors (which are perhaps to be localized within layer IV of the primary visual cortex, see, for instance, Usrey et al. (2003)) and two layers, Stage 1 (denoted as $S1$) and Stage 2 (denoted as $S2$) whose neurons are thought to be also localized in the same primary visual cortex area, presumably within layers II and III.
@@
Such an image is used as afferent input to the feature extraction neurons in $S1$ and first increases their “membrane potential” $\boldsymbol{a}^{(S1)}\in\mathbb{R}^{C^{(S1)}\times H\times W}$ and potentially triggers their binary activation state $\boldsymbol{y}^{(S1)}\in\{0,1\}^{C^{(S1)}\times H\times W}$ (green box in Figure 2). The membrane potential $\boldsymbol{a}^{(S1)}$ is calculated by applying a convolutional operation (LeCun et al., 1989) (yellow box $S1$ in Figure 2):
@@
with the same bias parameter $b^{(S1)}$ for all neurons acting as a firing threshold (cells fire if the membrane potential is above this threshold).
@@
Stage $S2$ builds net fragments by allowing active neurons to support laterally connected neurons over T time steps (depicted as recurrent connection in Figure 2). A sample image is introduced into the system at time step $t=0$ to activate neurons and remains unchanged until time step $t=T$. Throughout this period, neuronal activity undergoes sparsification due to growing inhibition. Through the sparsification process, the system forms an object representation by leaving active a set of neurons each of which is supported by active neighbors, that is, a net.
Short-Range Lateral Connections.
Although typical convolutional neural networks (Fukushima, 1980; Waibel et al., 1987; LeCun et al., 1989) apply convolution operations to establish feed-forward connections between different layers, we employ them here to implement short-range lateral connections within the same layer, their symmetry permitting a pattern to manifest itself at any spatial location based on corresponding net fragments. Given lateral connections, the neural activity of the previous time step $t-1$ is accessed to calculate a neuron’s activity at time $t$. In contrast to recurrent networks like LSTMs (Hochreiter and Schmidhuber, 1997), where a new token is input at each time step, we iterate over the same input while updating the internal state.
The membrane potentials $\boldsymbol{a}^{(S2)}_{c,j}$ of the second stage neurons get excited by the (constant) signals of the previous stage $S1$ and iteratively by lateral excitation from within $S2$. Before digitization, they are modified by saturation, normalization, and inhibition (yellow box $S2$ in Figure 2). Before those modifications, its raw form is computed as:
$\boldsymbol{a}^{\prime(S2)}_{c,j}[t]=\underbrace{\left(\sum_{k\in C^{(S1)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(F)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S1)}\right)}_{\textit{Forward Connection}}+\overbrace{\left(\sum_{k\in C^{(S2)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(L)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S2)}[t-1]\right)}^{\textit{Lateral Connection}}.$ (3)
@@
The forward connections are denoted as $\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)}\times C^{(S1)}\times h^{(W2)}\times w^{(W2)}}$ and the lateral connections as $\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)}\times C^{(S2)}\times h^{(W2)}\times w^{(W2)}}$, where $C^{(S2)}$ is the number of channels in $S2$, including the set of all feature channels and its alternative channels, i.e. $C^{(S2)}=n_{a}\cdot C^{(S1)}$ where $n_{a}$ denotes the number of alternative channels per feature type (the alternative channels handle the alternative neurons that are required to prevent cross-talk between fragments, see Apendix A). In the case of the forward connections, the kernel size determines which of the neurons $\boldsymbol{y}^{(S1)}$ in stage 1 can be connected to a neuron $\boldsymbol{y}^{(S2)}$ in stage 2 (red box in Figure 2), while the kernel size of the lateral connections determines which neurons can be connected within stage 2. The kernel $\boldsymbol{W}^{(F)}$ is initialized such that the sets of equivalent neurons in $S2$ copy the activity (and with that the feature type) of the corresponding neuron in $S1$, and $\boldsymbol{W}^{(L)}$ is initialized with zeros except for self-coupling of neurons, with value 1 (details in Appendix B). The weights are updated using Hebbian Learning (details in Appendix C).
For the first time step $t=0$, when $\boldsymbol{y}_{i}^{(S2)}$ is undefined, we initialize the $S2$ neurons with zeros, i.e., the neurons are silent and do not provide lateral support:
@@
This entire process is visually depicted in the upper part (A) of Figure 2. For simplicity in our algorithm, and since we use the same kernel size for the forward and lateral connections, we stack the two kernels $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$ as $\boldsymbol{W}^{(S2)}=(\boldsymbol{W}^{(F)};\boldsymbol{W}^{(L)})$ as well as the inputs $\boldsymbol{y}_{k,i}^{(S1)}$ and $\boldsymbol{y}_{k,i}^{(S2)}[t-1]$ as shown in Figure 2(B), leading to the equivalent formulation:
@@
The raw neural excitation $\boldsymbol{a}^{\prime}$ computed in eq. 3 (or eq. 5) is to be subjected to saturation, inhibition, and normalization so it can be binarized using a pre-defined threshold. These steps of processing the raw membrane potential $\boldsymbol{a}^{\prime}$ of neurons proved necessary in our simulation experiments to achieve the qualitative function we aimed for. The details to obtain the normalized activations $\boldsymbol{a}^{\text{norm}}$ in the range $(0,1)$ are described in Appendix D. The differences between activation levels of neurons are accentuated by setting
@@
The coefficient $\gamma$ regulates the inhibition strength and increases with the time-steps of the inner loop according to $\gamma=1.2+0.2t$, making it harder for neurons to fire (as $\boldsymbol{a}^{\text{norm }(S2)}\leq 1$). The membrane potential is higher for neurons that are part of coherent fragments (valid patterns) as they receive lateral support, while invalid activations are less supported. While the inhibition coefficient increases over time, many neurons eventually get suppressed, leaving only those active that support each other and form a consistent net. Finally, the membrane potential is converted to a binary activation, neurons becoming active when their inhibited membrane potential is above the activation bias $b^{(S2)}$, which serves to control the firing rate of neurons:
@@
The final output $\boldsymbol{y}^{(S2)}[t=T]$ is the lateral connection-supported neuronal activation in which, in contrast to the initial activation, unsupported activations are suppressed. To allow features to be part of multiple net fragments and represent various patterns, they are represented by a set of alternative neurons (details in Appendix A). All neurons within such sets are free to learn different lateral connections to other neurons. They are connected by distinct channels within the convolutional matrix and engage in competition, with only one alternative channel gaining activation at each spatial location while the remaining neurons are suppressed (set to zero). This selection process process is described in the Appendix E.
@@ 5 Experiments
We evaluate our framework by assessing the efficacy of net fragments to filter out random Gaussian noise, the capacity to reconstruct occluded objects, i.e., the efficacy of dealing with subtractive noise, and its generalization capacity by composing small net fragments into object structures not observed during training.
5.1 Dataset
The images are generated when required and their set may vary from one epoch to another. During each training cycle, we randomly generate $300$ image instances (each line about $5$ times). During evaluation, we sample each of the $59$ distinct lines exactly once. In contrast to training, the samples used for evaluation comprise local distortions in the form of additive Gaussian nose and missing line segments (subtractive noise), and we evaluate the net fragments’ capability to remove these.
We introduce additive Gaussian noise to the afferent input in stage $S2$, thereby simulating an image structure that has already survived the filtering of feature kernels in $S1$. To that end, we probabilistically flip neurons within each channel with a probability of up to $20\%$. Since we use four feature channels in our experiments, this corresponds to a probability of $1-(1-0.2)^{4}=59\%$ that a neuron is flipped at any spatial location. To evaluate subtractive noise, we create discontinuous lines by deleting a line segment in the middle.
To demonstrate the generalization capability of DNA areas, we generate validation samples that differ from the ones used during training. Specifically, we produce kinked lines by generating four random points within the image and draw straight lines between these points.
@@ 6 Results
In the following, we show that the proposed net fragments can suppress Gaussian noise in Section 6.1 and can reconstruct removed pixels, see Section 6.2. A description of the feature extraction mechanism to obtain $\boldsymbol{y}^{(S1)}$ is in Appendix F, and details about the used parameters are in Appendix G. Additionally, Appendix H contains visualizations of the average support strength received by cells, and Appendix I a quantitative evaluation of suppressed noise.
@@ 6.1 Filtering Gaussian Noise
Figure 3: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to suppression of neurons not supported by lateral connections of net fragments. In each image, the background is shown in black, and the four different features in distinct colors. The plots are created using a bias $b^{(S2)}=0.5$ and an inhibition coefficient $\gamma=1.2+0.2t$.
The illustration in Figure 3 visually demonstrates the system’s effective capability of filtering out isolated noise or small clusters thereof. Noise of varying degrees (from $0\%$ to $59\%$ noise probability at each spatial location - see Appendix 5.1) is introduced in an area’s afferent input, from which a significant portion is removed when constructing net fragments. While introducing up to $47.8\%$ noise, remarkably consistent outputs $\boldsymbol{y}^{(S2)}$ are generated. As additional noise is introduced, more undesired neurons receive support and persist in their activity. However, the overarching pattern remains distinguishable even when subjected to noise levels of up to $59\%$. In Appendix I, we quantitatively confirm the efficacy of noise filtering by measuring precision, recall, and noise filtration rate for various parameter settings, showing that lines remain well distinguishable for up to $59\%$ added noise.
@@
Figure 3 visualizes that neural activity simulating visual clutter without coherent structure is removed after processing for $5$ time-steps. Figure 4 depicts this systematic filtering over time. The additionally introduced clutter doesn’t activate overlapping net fragments and cannot withstand increasing inhibition. The figure illustrates a substantial noise reduction already at time step $t=0$ due to the filtering by matrix $\boldsymbol{W}^{(F)}$ and initial inhibition. In each subsequent time step, additional clutter is removed. This reduction is due to increased inhibition and falling support as a consequence of the drop-out of supporting neurons. We thus observe a chain reaction of gradual suppression of incoherent structure as well as the preservation of net structure on the basis of internal support.
6.2 Reconstructing Subtractive Noise
Figure 5: Partial occlusion as subtractive noise: Input images of diagonal lines with different numbers of missing pixels, denoted in the top left corner of the panels (left column), corresponding feature activations (middle column), and final output neuron activations. Activation bias $b^{(S2)}=0.5$, inhibition coefficient $\gamma=0.6+0.2t$.
If a substantial number of neurons activate and support other neurons within a net fragment, inactive neurons can receive significant support, leading them to switch on and encouraging figure reconstruction. Figure 5 shows how inactive neurons activate and thus demonstrates that net fragments can deal with subtractive noise (occluded patterns). This reconstruction works only reliably for up to $3$ missing pixels (see Appendix I). In some cases, as in this Figure, the line is fully reconstructed for up to $6$ removed pixels and partly reconstructed for $7$ pixels.

### v2

@@ 3 Computational Framework
Figure 2: The dynamics of a single area within the proposed CNA: The input image is fed into stage $S1$ to obtain feature activations. The feature activations are then processed by stage $S2$, together with activations from neurons within the same layer. In part A, two distinct matrices are used to model the forward connections from $S1$ to $S2$ and the recurrent connections within $S2$. In part B, the matrices and input signals are concatenated, resulting in one bigger weight matrix.
Our framework addresses the visual system, even though it could be applied to other modalities as well. In the primary visual cortex, images appear as two-dimensional arrays of activated neurons, and its neurons are sensitive to local texture elements within their receptive fields (Grill-Spector and Malach, 2004). To account for the homogeneity of image structure, it can be assumed that the connectivity has translational symmetry (‘convolutional’ structure) supporting equivariance, that is, the property that image shift gets reflected in activity shift. We use these insights to build a CNA with two stages, as shown in Figure 2(A). The CNA is based on an input layer representing observed images by simple feature detectors and two layers, Stage 1 (denoted as $S1$) and Stage 2 (denoted as $S2$).
@@
Such an image is used as afferent input to the feature extraction neurons in $S1$ and first increases their ‘membrane potential’ $\boldsymbol{a}^{(S1)}\in\mathbb{R}^{C^{(S1)}\times H\times W}$ and potentially triggers their binary activation state $\boldsymbol{y}^{(S1)}\in\{0,1\}^{C^{(S1)}\times H\times W}$ (green box in Figure 2). The membrane potential $\boldsymbol{a}^{(S1)}$ is calculated by applying a convolutional operation (LeCun et al., 1989) (yellow box $S1$ in Figure 2):
@@
with the same bias parameter $b^{(S1)}$ for all neurons acting as a firing threshold (neurons fire if the membrane potential is above this threshold).
@@
Stage $S2$ builds net fragments by allowing active neurons to support recurrently connected neurons over T time steps (depicted as recurrent connection in Figure 2). A sample image is introduced into the system at time step $t=0$ to activate neurons and remains unchanged until time step $t=T$. Throughout this period, neuronal activity undergoes sparsification due to growing inhibition (see below). Through the sparsification process, the system forms an object representation by leaving active a set of neurons each of which is supported by active neighbors, that is, a net.
Short-Range recurrent Connections.
Although typical convolutional neural networks (Fukushima, 1980; Waibel et al., 1987; LeCun et al., 1989) apply convolution operations to establish feedforward connections between different layers, we employ them here to implement short-range recurrent connections within the same layer, their symmetry permitting a pattern to manifest itself at any spatial location based on corresponding net fragments. Given recurrent connections, the neural activity of the previous time step $t-1$ is accessed to calculate a neuron’s activity at time $t$. In contrast to recurrent networks like LSTMs (Hochreiter and Schmidhuber, 1997), where a new token is input at each time step, we iterate over the same input while updating the internal state to project out a net representing a given sensory signal. Importantly, the internal state evolves continuously, even when processing the same input pattern, driven by dynamic changes in the support between neurons along recurrent connections.
The membrane potentials $\boldsymbol{a}^{(S2)}_{c,j}$ of the second stage neurons get excited by the (constant) signals of the previous stage $S1$ and iteratively by recurrent excitation from within $S2$. Before digitization, they are modified by saturation, normalization, and inhibition (yellow box $S2$ in Figure 2). Before those modifications, its raw form is computed as:
$\boldsymbol{a}^{\prime(S2)}_{c,j}[t]=\underbrace{\left(\sum_{k\in C^{(S1)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(F)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S1)}\right)}_{\textit{Forward Connection}}+\overbrace{\left(\sum_{k\in C^{(S2)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(L)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S2)}[t-1]\right)}^{\textit{Recurrent Connection}}.$ (3)
@@
The forward connections are denoted as $\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)}\times C^{(S1)}\times h^{(W2)}\times w^{(W2)}}$ and the recurrent connections as $\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)}\times C^{(S2)}\times h^{(W2)}\times w^{(W2)}}$, where $C^{(S2)}$ is the number of channels in $S2$, including the set of all feature channels and its alternative channels, i.e. $C^{(S2)}=n_{a}\cdot C^{(S1)}$ where $n_{a}$ denotes the number of alternative channels per feature type (the alternative channels handle the alternative neurons that are required to prevent cross-talk between fragments). In the case of the forward connections, the kernel size determines which of the neurons $\boldsymbol{y}^{(S1)}$ in stage 1 can be connected to a neuron $\boldsymbol{y}^{(S2)}$ in stage 2 (red box in Figure 2), while the kernel size of the recurrent connections determines which neurons can be connected within stage 2. The kernel $\boldsymbol{W}^{(F)}$ is initialized such that the sets of equivalent neurons in $S2$ copy the activity (and with that the feature type) of the corresponding neuron in $S1$, and $\boldsymbol{W}^{(L)}$ is initialized with zeros except for self-coupling of neurons, with value 1 (details in Appendix A). The weights are updated using Hebbian Learning (details in Appendix B).
For the first time step $t=0$, when $\boldsymbol{y}_{i}^{(S2)}$ is undefined, we initialize the $S2$ neurons with zeros, i.e., the neurons are silent and do not provide recurrent support:
@@
This entire process is visually depicted in the upper part (A) of Figure 2. For simplicity in our algorithm, and since we use the same kernel size for the forward and recurrent connections, we stack the two kernels $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$ as $\boldsymbol{W}^{(S2)}=(\boldsymbol{W}^{(F)};\boldsymbol{W}^{(L)})$ as well as the inputs $\boldsymbol{y}_{k,i}^{(S1)}$ and $\boldsymbol{y}_{k,i}^{(S2)}[t-1]$ as shown in Figure 2(B), leading to the equivalent formulation:
@@
The raw neural excitation $\boldsymbol{a}^{\prime}$ computed in eq. 3 (or eq. 5) is to be subjected to saturation, inhibition, and normalization so it can be binarized using a pre-defined threshold. These steps of processing the raw membrane potential $\boldsymbol{a}^{\prime}$ of neurons proved necessary in our simulation experiments to achieve the qualitative function we aimed for. The details to obtain the normalized activations $\boldsymbol{a}^{\text{norm}}$ in the range $(0,1)$ are described in Appendix C. The differences between activation levels of neurons are accentuated by setting
@@
The coefficient $\gamma$ regulates the inhibition strength and increases with the time-steps of the inner loop according to $\gamma=1.2+0.2t$, making it harder for neurons to fire (as $\boldsymbol{a}^{\text{norm }(S2)}\leq 1$). The membrane potential is higher for neurons that are part of coherent fragments (valid patterns) as they receive recurrent support, while invalid activations are less supported. While the inhibition coefficient increases over time, many neurons eventually get suppressed, leaving only those active that support each other and form a consistent net. Finally, the membrane potential is converted to a binary activation, neurons becoming active when their inhibited membrane potential is above the activation bias $b^{(S2)}$, which serves to control the firing rate of neurons:
@@
The final output $\boldsymbol{y}^{(S2)}[t=T]$ is the recurrent connection-supported neuronal activation in which, in contrast to the initial activation, unsupported activations are suppressed. Alternative neurons are connected by distinct channels within the convolutional matrix and engage in competition, with only one alternative channel gaining activation at each spatial location while the remaining neurons are suppressed (set to zero). This selection process process is described in the Appendix D.
Relation to Nets and Fragments.
Net fragments are learned through the strengthening of recurrent connections $\boldsymbol{W}^{(L)}$, facilitated by Hebbian plasticity. Neurons within a fragment are interconnected, providing mutual support that enhances their membrane potentials. By shifting $\boldsymbol{W}^{(L)}$ as convolutional kernels across the input, fragments can be detected at all spatial locations, resulting in numerous overlapping fragments. Coherent fragments are interconnected to form a larger structure, a net.
While multiple nets may exist, projecting out the most coherent one is crucial for robust object representation. This is achieved by progressively increasing inhibition over time. As inhibition intensifies, neurons with insufficient support (e.g., only a few instead of many neurons within one net fragment were triggered) are deactivated [^2] Neuron activations can be interpreted as hypotheses that require confirmation/support from other neurons with compatible hypotheses - otherwise, they are deactivated.. This deactivation can propagate, leading to the elimination of entire fragments and leaving a single coherent net to represent the input structure.
In our experiments (see below), we observe that, initially, neurons operate independently due to the absence of established recurrent connections, with a single neuron having a membrane potential of $1$ being sufficient for activation. However, after training, the average membrane potential rises to approximately $12$, reflecting robust support from neurons within a fragment. As a result, inconsistent neurons can no longer activate (and neurons with a membrane potential of $1$ are switched off), ensuring a more stable and coherent representation.
@@ 4 Experiments
To demonstrate the compositionality of net fragments, we train (updating $\boldsymbol{W}^{(L)}$ with Hebbian plasticity) the model on binary images containing straight-line patterns. For evaluation (after training and without updating any connections), we use images of kinked lines, digits, characters, and line drawings of objects such as houses and clocks. These complex patterns can be expressed as compositions of straight-line net fragments, even though the CNA was not explicitly trained on such patterns. This flexible compositional behavior contrasts strongly with associative memory (see Section 2.1). To highlight the differences between ANNs, we train and evaluate an autoencoder on the same dataset and compare the representations.
Besides compositionality, we introduce noise to examine the robustness and persistence of coherent structures with statistical relevance. Specifically, we assess the efficacy of filtering out random Gaussian noise and reconstructing partially occluded objects, i.e., the efficacy of dealing with subtractive noise. We introduce additive Gaussian noise to the afferent input in stage $S2$, thereby simulating an image structure that has already survived the filtering of feature kernels in $S1$ (thus increasing the difficulty, as the fragments must handle all the noise rather than relying on the feature extraction stage).
Further experimental details are provided in the Appendix: We describe the training and testing datasets in more detail in Appendix E, the feature extractor mechanism to obtain $\boldsymbol{y}^{(S1)}$ is described in Appendix F, the training details of our CNA model are in Appendix G, and the training details of the autoencoder in Appendix H.
@@ 5 Results
Figure 3: Net formed by activated neurons. The left of the image shows the neuronal activation $\boldsymbol{y}^{(S2)}$ when observing a line, and the right visualizes the net formed by these neurons. For convenience, the neurons are rearranged in a circle, and the connections of two randomly selected neurons, A and B, are visualized in colors (purple and red).
We empathize that this work focuses on learning a new neural code and, in the following, demonstrates it on a rather simple dataset.
Although our result figures display activated neurons, we would like to emphasize that these neurons are carriers of connections and form part of nets. To make this point for once, Figure 3 depicts this connectivity of activated neurons in $\boldsymbol{y}^{(S2)}$ (arranged as a circle for convenience).
5.1 Compositionality
Figure 4: The top of the figure illustrates the straight lines used during the training phase. Below, various samples from the evaluation datasets are shown, including kinked lines (rows $1$ and $2$), digits and characters (rows 3 and 4), and line drawings (rows $5$-$8$). For each row, two input examples, denoted as $\boldsymbol{x}_{1}$ and $\boldsymbol{x}_{2}$, are displayed alongside their respective feature activations, $\boldsymbol{y}^{(S1)}$, and neuronal activations, $\boldsymbol{y}^{(S2)}$, shown both without and with addition of Gaussian noise.
The model weights, $\boldsymbol{W}^{(S2)}$, are trained using straight lines (see training data at the top of Figure 4), allowing the model to capture the underlying structure of straight lines within net fragments. The model composes these fragments dynamically into nets to represent much more complex structures (see validation data in rows 1-8 of Figure 4). For instance, despite not encountering these images during training, the CNA model can represent kinked lines (rows $1$ and $2$), digits and characters (rows $3$ and $4$), and line drawings (rows $5$-$8$).
The motivation for why we build net fragments ($\boldsymbol{y}^{(S2)}$) and not just use the feature activation $\boldsymbol{y}^{(S1)}$ is demonstrated by introducing additional clutter: For each sample, we generate an input representation both with and without Gaussian noise. This noise corresponds to invalid patterns not observed during training, which lack sufficient recurrent support, leading to the deactivation of the corresponding neurons. Thus, the CNA model only allows valid sub-patterns to persist and be composed into novel structures, while invalid patterns, such as noise, are suppressed.
@@ 5.2 Filtering Gaussian Noise
Figure 5: Noise-corrupted feature activations $\boldsymbol{y}^{(S1)}$ alongside the corresponding output $\boldsymbol{y}^{(S2)}$, in which noise is reduced due to suppression of neurons not supported by recurrent connections of net fragments. In each image, the background is shown in black, and the four different features in distinct colors. The plots are created using a bias $b^{(S2)}=0.5$ and an inhibition coefficient $\gamma=1.2+0.2t$.
The illustration in Figure 5 visually demonstrates the system’s effective capability of filtering out isolated noise or small clusters thereof. Noise of varying degrees (from $0\%$ to $59\%$ noise at each spatial location - see Appendix E) is introduced in an area’s afferent input, from which a significant portion is removed when constructing net fragments. While introducing up to $47.8\%$ noise, remarkably consistent outputs $\boldsymbol{y}^{(S2)}$ are generated. As additional noise is introduced, more undesired neurons receive support and persist in their activity. However, the overarching pattern remains distinguishable even when subjected to noise levels of up to $59\%$. In Appendix I, we quantitatively confirm the efficacy of noise filtering by measuring precision, recall, and noise filtration rate for various parameter settings, showing that lines remain well distinguishable for up to $59\%$ added noise.
@@
Figure 6 depicts this systematic filtering over time. The additionally introduced clutter doesn’t activate overlapping net fragments and cannot persist over time. The figure illustrates a substantial noise reduction already at time step $t=0$ due to the filtering by matrix $\boldsymbol{W}^{(F)}$ and initial inhibition. In each subsequent time step, additional clutter is removed. This reduction is due to increased inhibition and falling support as a consequence of the drop-out of supporting neurons, demonstrating how overlapping fragments build a coherent net of mutually supporting neurons.
5.3 Reconstructing Figures
Figure 7: Part A: Input images of diagonal lines with different numbers of missing pixels, denoted in the top left corner of the panels (left column), corresponding feature activations (middle column), and final output neuron activations. Activation bias $b^{(S2)}=0.5$, inhibition coefficient $\gamma=0.6+0.2t$. Part B: Diagonal line with $6$ missing pixels in the center (shown on the left), the corresponding output activity (in the middle), and the membrane potential $\boldsymbol{a}^{(S2)}$displaying the ‘activation probabilities’ (on the right) after the first time step.
If a substantial number of neurons activate and support other neurons within a net, inactive neurons can receive significant support, leading them to switch on and encouraging figure reconstruction. Figure 7A shows how inactive neurons activate and thus demonstrates that nets can deal with occluded patterns. This reconstruction works only reliably for up to $3$ missing pixels (see Appendix I). In some cases, as in this Figure, the line is fully reconstructed for up to $6$ removed pixels and partly reconstructed for $7$ pixels.